OpenStudy (anonymous):

Three times a number plus 12 minus 5 times the same number is 22. What is the number?

OpenStudy (anonymous):

Start by assuming the number to be = x then keep building the equation as per the hints given

OpenStudy (anonymous):

I don't get it What is the answer

OpenStudy (anonymous):

Do u want just the answer or want to learn how to solve it???

OpenStudy (anonymous):

Answer

OpenStudy (carson889):

x = -5

OpenStudy (anonymous):

the answer is -5

OpenStudy (carson889):

Assuming the equation to be: 3x + 12 - (5x) = 22 3x - (5x) = 10 -2x = 10 x = -5

OpenStudy (anonymous):

THANK YOU EVERYBODY!